# EuleroDec: A Complex-Valued RVQ-VAE for Efficient and Robust Audio Coding

### v1

@@ Abstract
Audio codecs power discrete music generative modelling, music streaming, and immersive media by shrinking PCM audio to bandwidth-friendly bitrates. Recent works have gravitated towards processing in the spectral domain; however, spectrogram domains typically struggle with phase modeling, which is naturally complex-valued. Most frequency-domain neural codecs either disregard phase information or encode it as two separate real-valued channels, limiting spatial fidelity. This entails the need to introduce adversarial discriminators at the expense of convergence speed and training stability to compensate for the inadequate representation power of the audio signal. In this work we introduce an end-to-end complex-valued RVQ‑VAE audio codec that preserves magnitude-phase coupling across the entire analysis-quantization-synthesis pipeline and removes adversarial discriminators and diffusion post‑filters. Without GANs or diffusion, we match or surpass much longer-trained baselines in-domain and reach SOTA out-of-domain performance on phase coherence and waveform fidelity. Compared to standard baselines that train for hundreds of thousands of steps, our model, which reduces the training budget by an order of magnitude, is markedly more compute-efficient while preserving high perceptual quality.
@@ 1 Introduction
Spectral-domain audio codecs decompose a signal into its time-frequency representation, typically via short-time Fourier transform (STFT). Although the magnitude (or power) spectrum carries most perceptual content, the phase spectrum is equally essential for faithful reconstruction: improper or discarded phase information leads to the smearing of transients, ’phasiness’, and audible artifacts in the decoded signal. Many recent approaches, both in the waveform and spectral domains, employ multi-scale adversarial discriminators [defossez2022high] [kumar2023high] or score or flow-based post filters [welker2025flowdecflowbasedfullbandgeneral] [wu2024scoredec] to improve perceptual quality at the price of adversarial instability and slower convergence. An alternative line of work encodes complex spectra by splitting real-imaginary or magnitude–phase components and processing them independently [wu2025complexdec, du2024apcodecspectrumcodingbasedhighfidelityhighcompressionrate, ai2024apcodecneuralaudiocodec, du2023funcodecfundamentalreproducibleintegrable]. However, this factorization neglects intrinsic phase–magnitude coupling and leads to parameter redundancy. Complex-valued neural networks (CVNNs) have shown promise in speech enhancement [hu2020dccrndeepcomplexconvolution, halimeh2021complexvaluedspatialautoencodersmultichannel] and generative modeling [nakashika20_interspeech]. Yet existing codec-oriented approaches either operate partially in the real domain [Ru_2024], breaking phase coherence, or are restricted to music and a single bitrate. To date, to the best of our knowledge, no codec realizes a fully end-to-end complex-valued analysis-quantization-synthesis pipeline without adversarial or post-filtering components. In this work, we introduce the first fully end-to-end complex-valued neural audio codec for 24 kHz speech at 6 and 12 kbps. All stages of the pipeline - from waveform input to waveform reconstruction - are complex-valued and jointly optimized using Wirtinger calculus, without real-valued detours. The model integrates complex convolutions, attention, and normalization with 2×2 whitening, preserving the algebraic structure of the STFT and intrinsic amplitude–phase coupling. Without adversarial training or diffusion-based post-filters, the proposed approach achieves consistent in-domain and out-of-domain improvements at the target bitrates. Our contributions are as follows: (1) We propose the first end-to-end complex-valued VQ-VAE for audio, (2) our codec attains state-of-the-art quality without adversarial discriminators or score-based post-filters; (3) we reach and surpass state-of-the-art quality with fast, steady convergence and reduces training budget by $95\%$.

### v2

@@ Abstract
Audio codecs power discrete music generative modelling, music streaming and immersive media by shrinking PCM audio to bandwidth-friendly bit-rates. Recent works have gravitated towards processing in the spectral domain; however, spectrogram-domains typically struggle with phase modeling which is naturally complex-valued. Most frequency-domain neural codecs, either disregard phase information or encode it as two separate real-valued channels, limiting spatial fidelity. This entails the need to introduce adversarial discriminators at the expense of convergence speed and training stability to compensate for the inadequate representation power of the audio signal. In this work we introduce an end-to-end complex-valued RVQ‑VAE audio codec that preserves magnitude–phase coupling across the entire analysis–quantization–synthesis pipeline and removes adversarial discriminators and diffusion post‑filters. Without GANs or diffusion we match or surpass much longer-trained baselines in-domain and reach SOTA out-of-domain performance. Compared to standard baselines that train for hundreds of thousands of steps, our model reducing training budget by an order of magnitude is markedly more compute-efficient while preserving high perceptual quality.
@@ 1 Introduction
Spectral-domain audio codecs decompose a signal into its time-frequency representation, typically via short-time Fourier transform (STFT). Although the magnitude (or power) spectrum carries most perceptual content, the phase spectrum is equally essential for faithful reconstruction: improper or discarded phase information leads to the smearing of transients, ’phasiness’, and audible artifacts in the decoded signal. Many recent approaches, both in the waveform and spectral domains, employ multi-scale adversarial discriminators [defossez2022high] [kumar2023high] or score or flow-based post filters [welker2025flowdecflowbasedfullbandgeneral] [wu2024scoredec] [10890277] to improve perceptual quality at the price of adversarial instability and slower convergence. An alternative line of work encodes complex spectra by splitting real-imaginary or magnitude–phase components and processing them independently [wu2025complexdec, du2024apcodecspectrumcodingbasedhighfidelityhighcompressionrate, ai2024apcodecneuralaudiocodec, du2023funcodecfundamentalreproducibleintegrable]. However, this factorization neglects intrinsic phase–magnitude coupling and leads to parameter redundancy. Complex-valued neural networks (CVNNs) have shown promise in speech enhancement [hu2020dccrndeepcomplexconvolution, halimeh2021complexvaluedspatialautoencodersmultichannel] and generative modeling [nakashika20_interspeech]. Yet existing codec-oriented approaches either operate partially in the real domain [Ru_2024], breaking phase coherence, or are restricted to music and a single bitrate. To date, to the best of our knowledge, no codec realizes a fully end-to-end complex-valued analysis-quantization-synthesis pipeline without adversarial or post-filtering components. In this work, we introduce the first fully end-to-end complex-valued neural audio codec for 24 kHz speech at 6 and 12 kbps. All stages of the pipeline - from waveform input to waveform reconstruction - are complex-valued and jointly optimized using Wirtinger calculus, without real-valued detours. The model integrates complex convolutions, attention, and normalization with 2×2 whitening, preserving the algebraic structure of the STFT and intrinsic amplitude–phase coupling. Without adversarial training or diffusion-based post-filters, the proposed approach achieves consistent in-domain and out-of-domain improvements at the target bitrates. Our contributions are as follows: (1) We propose the first end-to-end complex-valued VQ-VAE for audio, (2) our codec attains state-of-the-art quality without adversarial discriminators or score-based post-filters; (3) we reach and surpass state-of-the-art quality with fast, steady convergence and reduces training budget by $95\%$.